OpenStudy (anonymous):

Which is not a factor of 48? 8 14 16 24

OpenStudy (anonymous):

14

OpenStudy (anonymous):

which on can't be divided by 48. so: 48/8=6 48/14=3.4 48/16=3 48/24=2

OpenStudy (anonymous):

Thank you guys so much!! Can you help with another one?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Which is a prime number? 15 27 41 57

OpenStudy (anonymous):

prime numbers are numbers that are above 1 and has no positive divisors other than 1 and itself

OpenStudy (anonymous):

it's 41. that's the only one with a factor of one and itself.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

you're welcome

OpenStudy (shamil98):

48/14=3.4 you can divide it by 14. 14 is not a factor of 48 because it does not produce a whole number. just a note.

OpenStudy (anonymous):

Thanks to every body who has helped me!!!